# The Nonlinear Filter Model of Stream Cipher Redivivus

### v6

@@ 4 Concrete Stream Cipher Proposals
We revisit the filter generator model where the state of an LFSR of length $L$ is filtered using a Boolean function. By $\mathcal{S}(L,m)$, we denote the class of stream ciphers obtained from an LFSR of length $L$ with a primitive connection polynomial, and with $f_{2m+1}$ as the filtering function. For the $\kappa$-bit security level, we assume that the stream cipher supports a $\kappa$-bit secret key and a $\kappa$-bit initialisation vector (IV), and so $L\geq 2\kappa$. Below we describe the construction of $\mathcal{S}(L,m)$ which requires specifying a few more parameters in addition to $L$ and $m$.
@@
${\sf nb}(u_{L-1},\ldots,u_{1},u_{0})$ $=$ $c_{L-1}u_{L-1}\oplus\cdots\oplus c_{1}u_{1}\oplus c_{0}u_{0},$
${\sf NS}(u_{L-1},\ldots,u_{1},u_{0})$ $=$ $({\sf nb}(u_{L-1},\ldots,u_{1},u_{0}),u_{L-1},\ldots,u_{1}).$
@@
The function $f_{2m+1}(W,\mathbf{X},\mathbf{Y})$, where $\mathbf{X}=(X_{1},\ldots,X_{m})$ and $\mathbf{Y}=(Y_{1},\ldots,Y_{m})$, is a function on $12m+$ 1 variables. This function will be used as the filtering function and applied to a subset of the bits of the $L$-bit state. We next describe the tap positions of the $L$-bit state which provides the input bits to $f_{2m+1}$.
Let $i_{1},\ldots,i_{m},j_{1},\ldots,j_{m}$ be integers satisfying the following condition.
$2\displaystyle 0=i_{1}<i_{2}<\cdots<i_{m-1}<i_{m}<\kappa\leq j_{1}<j_{2}<\cdots<j_{m-1}<j_{m}=2\kappa-$ 2 . (6)
@@
The function ${\sf proj}$ extracts the subset of $12m+$ 1 bits from the $L$-bit state to which the function $f_{2m+1}$ is to be applied. We define the composition $f_{2m+1}\circ{\sf proj}$ to be the following $L$-variable Boolean function.
@@
Note that the tap position for $W$ is $L-2\kappa$, the tap position for $X_{p}$ is $L-1-i_{p}$, and the tap position for $Y_{p^{\prime}}$ is $L-1-j_{p^{\prime}}$, where $1\leq p,p^{\prime}\leq m$. There is no tap position in the rightmost $L-2\kappa$ positions of the state; the leftmost position (i.e. position $L-1$) of the state is the tap position for $X_{1}$, all the tap positions for the variables in $\mathbf{X}$ are in the leftmost $\kappa$ positions of the state; all the tap positions for the variables in $\mathbf{Y}$ are in the next $\kappa$ positions of the state, and the position $L-2\kappa+1$ of the state is the tap position for $Y_{m}$. We have
@@
and ${\sf pos}$ is 0 at all the other $L-2m-1$ positions.
Note that the indexing of ${\sf pos}$ has $L-1$ as the leftmost position and 00 as the rightmost position. The string ${\sf pos}$ encodes the $12m+$ 1 tap positions of the state. Let ${\sf posX}$ be the leftmost $\kappa$-bit string of ${\sf pos}$, i.e. the segment of ${\sf pos}$ given by ${\sf pos}[L-1,\ldots,L-\kappa]$. The string ${\sf posX}$ encodes the tap positions of the variables in $\mathbf{X}$.
@@
The stream cipher uses a $\kappa$-bit key $(k_{\kappa-1},\ldots,k_{0})$ and a $\kappa$-bit initialisation vector $(v_{\kappa-1},\ldots,v_{0})$. Let $\mathbf{b}=(b_{L-2\kappa-1},\ldots,b_{0})$ be an $(L-2\kappa)$-bit string defined in the following manner: if $L-2\kappa$ is even, then $\mathbf{b}=(01)^{(L-2\kappa)/2}$, and if $L-2\kappa$ is odd, then $\mathbf{b}=1(01)^{(L-2\kappa-1)/2}$. The initial $L$-bit state $\mathbf{s}$ of the stream cipher is
@@
The initialisation round function is the map ${\sf IR}:\{0,1\}^{L}\rightarrow\{0,1\}^{L}$, where ${\sf IR}(u_{L-1},\ldots,u_{0})$ is defined in the following manner. Let $c={\sf nb}(u_{L-1},\ldots,u_{0})$ and $b=f_{2m+1}({\sf proj}(u_{L-1},\ldots,u_{0}))$. Then ${\sf IR}(u_{L-1},\ldots,u_{0})=(w_{L-1},\ldots,w_{0})$, where $w_{L-1}=c\oplus b$, and for $i=0,\ldots,L-2$,
$w_{i}$ $=$ $\left\{\begin{array}[]{ll}u_{i+1}\oplus b&\mbox{if }i\geq L-2\kappa\mbox{ and }i\equiv 0\bmod\mu,\\ u_{i+1}&\mbox{otherwise.}\end{array}\right.$
In the above, $\mu$ is a parameter of the initialisation round function. The round function feeds back the output of $f_{2m+1}\circ{\sf proj}$ to the state at $\lceil 2\kappa/\mu\rceil$ positions. This requires $\lceil 2\kappa/\mu\rceil$ XOR gates. We propose
$\mu$ $=$ $\lfloor\sqrt{2\kappa}\rfloor$ (12)
so that the feedback does not require too many XOR gates.
@@ 4.1 Effect of Shifts on Tap Positions
In our concrete proposals, the number of variables $n=2m+1$ of the filtering function is almost about half the length $L$ of the LFSR (see Table 1). In fact, the ratio $n/L$ in our proposals is much higher than similar ratios in all previous proposals. Using a high value of $n$ prevents certain well known attacks as we discuss later. On the other hand, the high (compared to previous proposals) value of $n/L$ ratio means that the tap positions are placed close together. Since $L-1$ bits of the next state of the LFSR are obtained by a shift of the previous state, two states of the LFSR which are close in time share a number of bits. Further, with a high value of $n/L$ ratio, two keystream bits arising from two such nearby states may depend on a number of common state bits. This creates the possibility that in the XOR of two such keystream bits, a number of terms involving the state bits cancel out. A bad choice of the tap positions can indeed lead to such an effect, as is explained in Section 4.4. Below we show that such cancellations do not occur in our proposals.
@@
When applied to the bits of the state, the inner product $\langle{\sf rev}(\mathbf{X}),\mathbf{Y}\rangle$ leads to $m$ quadratic terms involving the state bits. Two states which are separated by $d$ time periods share a number of bits. More precisely, the rightmost $L-d$ bits of the later state are the leftmost $L-d$ bits of the former state. This creates the possibility that some of the quadratic terms of the former state cancels with some of the quadratic terms of the later state. We show that due to our choice of the permutation $\pi$ as the bit reversal permutation ${\sf rev}$ and the choice of placing the tap positions for $\mathbf{X}$ to the left of the tap positions for $\mathbf{Y}$, such cancellations do not occur.
Let $\mathbf{s}^{(t)}$ be the state at time point $t$ as defined as in (14). The $p$-th quadratic term of $\mathbf{s}^{(t)}$ arising from $\langle{\sf rev}(\mathbf{X}),\mathbf{Y})$ is $s_{L+t-1-i_{p}}s_{L+t-1-j_{m+1-p}}$, where $1\leq p\leq m$. Let $t_{1}\geq 0$ and $d>0$ be integers, and $t_{2}=t_{1}+d$. Let $\mathbf{s}^{(t_{1})}$ and $\mathbf{s}^{(t_{2})}$ be the states at time points $t_{1}$ and $t_{2}$ respectively (obtained from (14) by substituting $t_{1}$ and $t_{2}$ for $t$). If there are integers $p$ and $p^{\prime}$ with $1\leq p,p^{\prime}\leq m$ such that the $p$-th quadratic term of $\mathbf{s}^{(t_{1})}$ is identically equal to the $p^{\prime}$-th quadratic term of $\mathbf{s}^{(t_{2})}$, then we say that a cancellation of quadratic terms occur between $\mathbf{s}^{(t_{1})}$ and $\mathbf{s}^{(t_{2})}$.
@@
No cancellation of quadratic terms occur between $\mathbf{s}^{(t_{1})}$ and $\mathbf{s}^{(t_{2})}$.
@@
The quadratic terms arising from $\mathbf{s}^{(t_{1})}$ and $\mathbf{s}^{(t_{2})}$ are
$s_{L+t_{1}-1-i_{p}}s_{L+t_{1}-1-j_{m+1-p}}$ and $s_{L+t_{2}-1-i_{p^{\prime}}}s_{L+t_{2}-1-j_{m+1-p^{\prime}}}$ (13)
respectively, for $1\leq p,p^{\prime}\leq m$. Cancellation of the $p$-th quadratic term of $\mathbf{s}^{(t_{1})}$ with the $p^{\prime}$-th quadratic term of $\mathbf{s}^{(t_{2})}$ occurs if $s_{L+t_{1}-1-i_{p}}s_{L+t_{1}-1-j_{m+1-p}}\mbox{ and }s_{L+t_{2}-1-i_{p^{\prime}}}s_{L+t_{2}-1-j_{m+1-p^{\prime}}}$ are identically equal. There are two cases to consider.
Case 1: This case corresponds to the possibility that $s_{L+t_{1}-1-i_{p}}$ and $s_{L+t_{2}-1-i_{p^{\prime}}}$ are identically equal and further $s_{L+t_{1}-1-j_{m+1-p}}$ and $s_{L+t_{2}-1-j_{m+1-p^{\prime}}}$ are identically equal. These two conditions are equivalent to $L+t_{1}-1-i_{p}=L+t_{2}-1-i_{p^{\prime}}$ and $L+t_{1}-1-j_{m+1-p}=L+t_{2}-1-j_{m+1-p^{\prime}}$, i.e. $i_{p^{\prime}}=i_{p}+d$ and $j_{m+1-p^{\prime}}=j_{m+1-p}+d$. Since $d>0$, we have $i_{p^{\prime}}>i_{p}$ and from the definition of $i_{1},\ldots,i_{m}$, we obtain $p^{\prime}>p$. Similarly, we have $j_{m+1-p^{\prime}}>j_{m+1-p}+d$, and from the definition of $j_{1},\ldots,j_{m}$, we obtain $m+1-p^{\prime}>m+1-p$, i.e. $p>p^{\prime}$, which is a contradiction to $p^{\prime}>p$.
Case 2: This case corresponds to the possibility that $s_{L+t_{1}-1-i_{p}}$ and $s_{L+t_{2}-1-j_{m+1-p^{\prime}}}$ are identically equal and further $s_{L+t_{2}-1-i_{p^{\prime}}}$ and $s_{L+t_{1}-1-j_{m+1-p}}$ are identically equal. These two conditions are equivalent to $L+t_{1}-1-i_{p}=L+t_{2}-1-j_{m+1-p^{\prime}}$ and $L+t_{2}-1-i_{p^{\prime}}=L+t_{1}-1-j_{m+1-p}$, i.e. $j_{m+1-p^{\prime}}=i_{p}+d$ and $i_{p^{\prime}}=j_{m+1-p}+d$. Since $d>0$, we obtain $i_{p^{\prime}}>j_{m+1-p}$, but by definition $i_{p^{\prime}}<j_{m+1-p}$ (see (6)), and so we obtain a contradiction.
Suppose two keystream bits are XORed. Proposition 5 assures us that the quadratic terms arising in the expressions for these keystream bits do not cancel out with each other. During the keystream generation, suppose that at some point the state is $\mathbf{s}^{(0)}=(s_{L-1},\ldots,s_{0})$. For $t\geq 1$, define $s_{L+t-1}={\sf nb}(s_{L+t-2},\ldots,s_{t-1})$, and
$\mathbf{s}^{(t)}$ $=$ $(s_{L+t-1},s_{L+t-2},\ldots,s_{t}).$ (14)
@@
To generate keystream bits, the function $f\circ{\sf proj}$ is applied to $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$. This in turn requires applying the majority function to a subset of the bits of $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$. More precisely, the functions calls ${\sf Maj}(s_{L-1-i_{1}},\ldots,s_{L-1-i_{m}})$ and ${\sf Maj}(s_{L+t-1-i_{1}},\ldots,s_{L+t-1-i_{m}})$ occur corresponding to $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$ respectively. If $t<\kappa$, then there is a possibility that the inputs to the two calls to majority have an overlap, i.e. some of the inputs are common to both the majority calls. Since majority is a symmetric function, the actual positions where the common inputs occur do not matter. For $t\geq 1$, let
$\nu_{t}$ $=$ #($\#\left(\{L-1-i_{1},\ldots,L-1-i_{m}\}\cap\{L+t-1-i_{1},\ldots,L+t-1-i_{m}\}\right).$ (15)
@@
where $\wedge$ denotes the bitwise AND operation, and tmuch-greater-$\gg t$ denotes right shift by $t$ places. Define

### v7

@@ 4 Concrete Stream Cipher Proposals
We revisit the filter generator model where the state of an LFSR of length $L$ is filtered using a Boolean function. We refer to [38] for the general theory of LFSRs.
In Figure 1 we provide a schematic diagram of the nonlinear filter model of stream cipher. By $\mathcal{S}(L,m)$, we denote the class of stream ciphers obtained from an LFSR of length $L$ with a primitive connection polynomial, and with $f_{2m+1}$ as the filtering function. For the $\kappa$-bit security level, we assume that the stream cipher supports a $\kappa$-bit secret key and a $\kappa$-bit initialisation vector (IV), and so $L\geq 2\kappa$. Below we describe the construction of $\mathcal{S}(L,m)$ which requires specifying a few more parameters in addition to $L$ and $m$.
[…]
Figure 1: Schematic diagram of keystream generation by the nonlinear filter model of stream cipher. Here $r$ is the number of tap positions required to obtain the next bit of the LFSR.
@@
${\sf nb}(u_{L-1},\ldots,u_{1},u_{0})$ $=$ $c_{L-1}u_{L-1}\oplus\cdots\oplus c_{1}u_{1}\oplus c_{0}u_{0},$ (7)
${\sf NS}(u_{L-1},\ldots,u_{1},u_{0})$ $=$ $({\sf nb}(u_{L-1},\ldots,u_{1},u_{0}),u_{L-1},\ldots,u_{1}).$ (8)
It is easy to note that the map $\sf NS$} is efficiently invertible. We define the vector $\mathbf{c}$ to be the following.
$\mathbf{c}$ $=$ $(c_{L-1},\ldots,c_{1},c_{0}).$ (9)
The positions of $\mathbf{c}$ where the value is 1 are the positions of the LFSR that are tapped to obtain the next bit of the LFSR.
@@
The function $f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ (given by (6)), where $\mathbf{X}=(X_{1},\ldots,X_{m})$ and $\mathbf{Y}=(Y_{1},\ldots,Y_{m})$, is a function on $2m+1$ variables. This function will be used as the filtering function and applied to a subset of the bits of the $L$-bit state. We next describe the tap positions of the $L$-bit state which provides the input bits to $f_{2m+1}$.
In the concrete stream cipher proposals we put forward, we have $2m+1$ to be less than the security parameter $\kappa$. Hence, in what follows we assume $2m+1<\kappa$. Let $i_{1},\ldots,i_{m},j_{1},\ldots,j_{m}$ be integers satisfying the following condition.
$0=i_{1}<i_{2}<\cdots<i_{m-1}<i_{m}<\kappa\leq j_{1}<j_{2}<\cdots<j_{m-1}<j_{m}=2\kappa-2.$ (10)
@@
The function $\sf proj$} extracts the subset of $2m+1$ bits from the $L$-bit state to which the function $f_{2m+1}$ is to be applied. We define the composition $f_{2m+1}\circ{\sf proj}$ to be the following $L$-variable Boolean function.
@@
Note that the tap position for $W$ is $L-2\kappa$, the tap position for $X_{p}$ is $L-1-i_{p}$, and the tap position for $Y_{p^{\prime}}$ is $L-1-j_{p^{\prime}}$, where $1\leq p,p^{\prime}\leq m$. There is no tap position in the rightmost $L-2\kappa$ positions of the state; the leftmost position (i.e. position L1L-1) of the state is the tap position for $X_{1}$, all the tap positions for the variables in $\mathbf{X}$ are in the leftmost $\kappa$ positions of the state; all the tap positions for the variables in $\mathbf{Y}$ are in the next $\kappa$ positions of the state, and the position $L-2\kappa+1$ of the state is the tap position for $Y_{m}$. We have
@@
and $\sf pos$} is 0 at all the other L2m1L-2m-1 positions.
Note that the indexing of $\sf pos$} has L1L-1 as the leftmost position and 0 as the rightmost position. The string $\sf pos$} encodes the $2m+1$ tap positions of the state. Let $\sf posX$} be the leftmost $\kappa$-bit string of $\sf pos$}, i.e. the segment of $\sf pos$} given by ${\sf pos}[L-1,\ldots,L-\kappa]$. The string $\sf posX$} encodes the tap positions of the variables in $\mathbf{X}$.
@@
The stream cipher uses a $\kappa$-bit key $(k_{\kappa-1},\ldots,k_{0})$ and a $\kappa$-bit initialisation vector $(v_{\kappa-1},\ldots,v_{0})$. The state of the stream cipher is the state of the LFSR which is an $L$-bit string. In the initialisation phase, the concatenation of the key and the IV is padded to obtain an $L$-bit string which is loaded into the LFSR. Then the stream cipher is clocked for $2\kappa$ rounds. In these rounds the keystream bit that is generated is not produced as output. Instead it is fed back into the state of the LFSR. The goal of the initialisation phase is to obtain a good mix of the key and the IV bits, i.e. at the end of the initialisation phase each bit of the state is a “complex” nonlinear function of the key and the IV bits. We design the round function of the initialisaton phase to be an invertible map so that the entire initialisation phase is invertible. This ensures that the entropy of the key remains unchanged during the initialisation phase. Below we provide the details of the initialisation phase.
First we define the initialisation round function. To do this we need an $L$-bit string $\mathbf{d}=(d_{L-1},\ldots,d_{0})$. We put the following conditions on $\mathbf{d}$.
$\left.\begin{array}[]{rcll}d_{L-1}&=&1,&\\ d_{i}&=&0&\mbox{for }i=L-2\kappa-1,\ldots,0,\\ d_{i}&=&0&\mbox{if either }c_{i+1}=1\mbox{ or }{\sf pos}[i+1]=1,\mbox{ for }i=L-2,\ldots,L-2\kappa.\end{array}\right\}$ (17)
Given the string $\mathbf{d}$ we define the initialisation round function ${\sf IR}:\{0,1\}^{L}\rightarrow\{0,1\}^{L}$ as follows. For $\mathbf{u}=(u_{L-1},\ldots,u_{0})$, we define ${\sf IR}(\mathbf{u})$ to be the $L$-bit string $\mathbf{w}=(w_{L-1},\ldots,w_{0})$, where
$(w_{L-1},w_{L-2},\ldots,w_{0})$ $=$ $({\sf nb}(\mathbf{u}),u_{L-1},\ldots,u_{1})\oplus(d_{L-1}\mathfrak{b},d_{L-2}\mathfrak{b},\ldots,d_{0}\mathfrak{b}),$ (18)
with $\mathfrak{b}=f_{2m+1}({\sf proj}(u_{L-1},\ldots,u_{0}))$. Basically, on input $\mathbf{u}$ the function $\sf IR$} clocks the LFSR once and feeds back the keystream bit $\mathfrak{b}$ obtained from $\mathbf{u}$ to the positions of the state corresponding to the positions where the string $\mathbf{d}$ has the value 1. In (17), the first condition says that feedback is always provided to the leftmost bit of the state, the second condition says that no feedback is provided to the rightmost $L-2\kappa$ bits of the state, and the third condition says that if the $i$-th position is tapped either for the next state bit of the LFSR, or for the input to the function $f_{2m+1}$, then no feedback is provided to the (i1)(i-1)-st position of the state. This last condition on $\mathbf{d}$ is required in the next result which shows that the function $\sf IR$} is invertible.
Proposition 5.
The function $\sf IR$} defined in (18) is invertible.
Proof.
Let $\mathbf{w}={\sf IR}(\mathbf{u})$. We argue that the string $\mathbf{w}$ uniquely determines the string $\mathbf{u}$. From the conditions on $\mathbf{d}$ given in (17), we have $d_{L-1}=1$ and so from (18) we have $w_{L-1}={\sf nb}(\mathbf{b})\oplus\mathfrak{b}=c_{L-1}u_{L-1}\oplus\cdots\oplus c_{1}u_{1}\oplus c_{0}u_{0}\oplus\mathfrak{b}=c_{L-1}u_{L-1}\oplus\cdots\oplus c_{1}u_{1}\oplus u_{0}\oplus\mathfrak{b}$, since $c_{0}=1$ due to the connection polynomial of the LFSR being primitive. So $u_{0}=w_{L-1}\oplus c_{L-1}u_{L-1}\oplus\cdots\oplus c_{1}u_{1}\oplus\mathfrak{b}=w_{L-1}\oplus\mathfrak{c}\oplus\mathfrak{b}$, where $\mathfrak{c}=c_{L-1}u_{L-1}\oplus\cdots\oplus c_{1}u_{1}$. Further, for $i=L-1,\ldots,1$, $u_{i}=w_{i-1}\oplus d_{i-1}\mathfrak{b}$ and consequently, $u_{i}=w_{i-1}$ whenever $d_{i-1}=0$. From the conditions on $\mathbf{d}$ given in (17), $d_{i-1}=0$ if the $i$-th position is tapped either for computing ${\sf nb}(\mathbf{u})$, or for the input to the function $f_{2m+1}$. In other words, if $u_{i}$ is involved in either the computation of ${\sf nb}(\mathbf{u})$, or the computation of $\mathfrak{b}$, then $u_{i}$ is simply equal to $w_{i-1}$. So given $\mathbf{w}$, both the bits $\mathfrak{b}$ and $\mathfrak{c}$ are uniquely defined. As a result, $u_{0}=w_{L-1}\oplus\mathfrak{c}\oplus\mathfrak{b}$ is also uniquely defined, and so is $u_{i}=w_{i-1}\oplus d_{i-1}\mathfrak{b}$, for $i=L-1,\ldots,1$ and $d_{i-1}=1$.
While Proposition 5 states that $\sf IR$} is invertible, the proof actually shows that given $\mathbf{w}={\sf IR}(\mathbf{u})$ it is easy to obtain $\mathbf{u}$, i.e. $\sf IR$} is efficiently invertible.
Intuitively, feeding back the nonlinear bit $\mathfrak{b}$ to a large number of positions (subject to the constraints which ensure invertibility of the round function) potentially leads to a more complex and nonlinear state. On the other hand, the implementation of (18) requires a XOR gate for each $i$ such that $d_{i}=1$, for a total of ${\sf wt}(\mathbf{d})$ XOR gates. So too many feedback positions increases the gate count. As a compromise between these two considerations we propose that the weight of $\mathbf{d}$ be taken to be $\mu=\lfloor\sqrt{2\kappa}\rfloor$ and the positions of the 1’s in $\mathbf{d}$ are more or less equi-spaced so that the difference between two 1-positions is also about $\mu$. Later for the concrete stream cipher proposals, we provide the corresponding $\mathbf{d}$’s.
Next we show how $\sf IR$} is used to perform the initialisation phase. Let $(b_{L-2\kappa-1},\ldots,b_{0})$ be a bit string of length $L-2\kappa$ defined as follows: $(b_{L-2\kappa-1},\ldots,b_{0})$ is equal to $(01)^{(L-2\kappa)/2}$ if $L-2\kappa$ is even and $(b_{L-2\kappa-1},\ldots,b_{0})$ is equal to $1(01)^{(L-2\kappa-1)/2}$ if $L-2\kappa$ is odd. The string $(b_{L-2\kappa-1},\ldots,b_{0})$ is used to pad the concatenation of the key and IV to obtain a string of length $L$. The initial $L$-bit state $\mathbf{s}$ of the LFSR is
@@
Our choice of the padding string is somewhat arbitrary (other than being an almost balanced string). In particular, we do not know whether there is any security implication of using some other string for padding.
@@
Suppose that the state of the stream cipher at the end of the initialisation phase is $\mathbf{s}^{(0)}=(s_{L-1},\ldots,s_{0})$. For $t\geq 1$, define $s_{L+t-1}={\sf nb}(s_{L+t-2},\ldots,s_{t-1})$, and
$\mathbf{s}^{(t)}$ $=$ $(s_{L+t-1},s_{L+t-2},\ldots,s_{t}).$ (20)
Then $z_{t}=f_{2m+1}({\sf proj}(\mathbf{s}^{(t)}))$.
Length of keystream.
We stipulate that from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. To the best of our understanding, this is sufficient for all practical purposes.
@@ 4.1 Effect of Shifts on Tap Positions
In our concrete proposals, the number of variables $n=2m+1$ of the filtering function is almost about half the length $L$ of the LFSR (see Table 1). In fact, the ratio $n/L$ in our proposals is much higher than similar ratios in all previous proposals. Using a high value of $n$ prevents certain well known attacks as we discuss later. On the other hand, the high (compared to previous proposals) value of $n/L$ ratio means that the tap positions are placed close together. Since L1L-1 bits of the next state of the LFSR are obtained by a shift of the previous state, two states of the LFSR which are close in time share a number of bits. Further, with a high value of $n/L$ ratio, two keystream bits arising from two such nearby states may depend on a number of common state bits. This creates the possibility that in the XOR of two such keystream bits, a number of terms involving the state bits cancel out. A bad choice of the tap positions can indeed lead to such an effect, as is explained in Section 4.4. Below we show that such cancellations do not occur in our proposals.
@@
When applied to the bits of the state, the inner product $\langle{\sf rev}(\mathbf{X}),\mathbf{Y}\rangle$ leads to $m$ quadratic terms involving the state bits. Two states which are separated by $d$ time periods share a number of bits. More precisely, the rightmost LdL-d bits of the later state are the leftmost LdL-d bits of the former state. This creates the possibility that some of the quadratic terms of the former state cancels with some of the quadratic terms of the later state. We show that due to our choice of the permutation $\pi$ as the bit reversal permutation $\sf rev$} and the choice of placing the tap positions for $\mathbf{X}$ to the left of the tap positions for $\mathbf{Y}$, such cancellations do not occur.
Let $\mathbf{s}^{(t)}$ be the state at time point $t$ as defined as in (20). The $p$-th quadratic term of $\mathbf{s}^{(t)}$ arising from $\langle{\sf rev}(\mathbf{X}),\mathbf{Y})$ is $s_{L+t-1-i_{p}}s_{L+t-1-j_{m+1-p}}$, where $1\leq p\leq m$. Let
${\sf qterm}(t,p)$ $=$ $\{L+t-1-i_{p},L+t-1-j_{m+1-p}\},$ (21)
i.e. ${\sf qterm}(t,p)$ is the 2-element set of indices which define the $p$-th quadratic term of $s^{(t)}$.
Let $t_{1}\geq 0$ and $d>0$ be integers, and $t_{2}=t_{1}+d$. Let $\mathbf{s}^{(t_{1})}$ and $\mathbf{s}^{(t_{2})}$ be the states at time points $t_{1}$ and $t_{2}$ respectively (obtained from (20) by substituting $t_{1}$ and $t_{2}$ for $t$). If there are integers $p$ and $p^{\prime}$ with $1\leq p,p^{\prime}\leq m$ such that ${\sf qterm}(t_{1},p)$ is equal to ${\sf qterm}(t_{2},p^{\prime})$, then the $p$-th quadratic term of $\mathbf{s}^{(t_{1})}$ is identically equal to the $p^{\prime}$-th quadratic term of $\mathbf{s}^{(t_{2})}$. In this case a cancellation of quadratic terms occur between $\mathbf{s}^{(t_{1})}$ and $\mathbf{s}^{(t_{2})}$. The following result shows that such cancellations do not occur.
@@
For $0\leq t_{1}<t_{2}$ and $1\leq p,p^{\prime}\leq m$, ${\sf qterm}(t_{1},p)\neq{\sf qterm}(t_{2},p^{\prime})$.
@@
We have ${\sf qterm}(t_{1},p)=\{L+t_{1}-1-i_{p},L+t_{1}-1-j_{m+1-p}\}$ and ${\sf qterm}(t_{2},p^{\prime})=\{L+t_{2}-1-i_{p^{\prime}},L+t_{2}-1-j_{m+1-p^{\prime}}\}$. There are two cases to consider.
Case 1: This case corresponds to the possibility $L+t_{1}-1-i_{p}=L+t_{2}-1-i_{p^{\prime}}$ and $L+t_{1}-1-j_{m+1-p}=s_{L+t_{2}-1-j_{m+1-p^{\prime}}}$. These two conditions are equivalent to $L+t_{1}-1-i_{p}=L+t_{2}-1-i_{p^{\prime}}$ and $L+t_{1}-1-j_{m+1-p}=L+t_{2}-1-j_{m+1-p^{\prime}}$, i.e. $i_{p^{\prime}}=i_{p}+d$ and $j_{m+1-p^{\prime}}=j_{m+1-p}+d$. Since $d>0$, we have $i_{p^{\prime}}>i_{p}$ and from the definition of $i_{1},\ldots,i_{m}$, we obtain $p^{\prime}>p$. Similarly, we have $j_{m+1-p^{\prime}}>j_{m+1-p}+d$, and from the definition of $j_{1},\ldots,j_{m}$, we obtain $m+1-p^{\prime}>m+1-p$, i.e. $p>p^{\prime}$, which is a contradiction to $p^{\prime}>p$.
Case 2: This case corresponds to the possibility that $L+t_{1}-1-i_{p}=L+t_{2}-1-j_{m+1-p^{\prime}}$ and $L+t_{2}-1-i_{p^{\prime}}=s_{L+t_{1}-1-j_{m+1-p}}$. These two conditions are equivalent to $L+t_{1}-1-i_{p}=L+t_{2}-1-j_{m+1-p^{\prime}}$ and $L+t_{2}-1-i_{p^{\prime}}=L+t_{1}-1-j_{m+1-p}$, i.e. $j_{m+1-p^{\prime}}=i_{p}+d$ and $i_{p^{\prime}}=j_{m+1-p}+d$. Since $d>0$, we obtain $i_{p^{\prime}}>j_{m+1-p}$, but by definition $i_{p^{\prime}}<j_{m+1-p}$ (see (10)), and so we obtain a contradiction.
@@
While defining $s^{(t)}$ in (20), we considered $s^{(0)}$ to be the state after the initialisation phase. More generally, by shifting the time steps, we may consider $s^{(0)}$ to be the state at any point of time and $s^{(t)}$ to be the state obtained from $s^{(0)}$ after $t$ time steps. The function $f\circ{\sf proj}$ is applied to both the states $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$. This in turn requires applying the majority function to a subset of the bits of $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$. More precisely, the functions calls ${\sf Maj}(s_{L-1-i_{1}},\ldots,s_{L-1-i_{m}})$ and ${\sf Maj}(s_{L+t-1-i_{1}},\ldots,s_{L+t-1-i_{m}})$ occur corresponding to $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$ respectively. If $t<\kappa$, then there is a possibility that the inputs to the two calls to majority have an overlap, i.e. some of the inputs are common to both the majority calls. Since majority is a symmetric function, the actual positions where the common inputs occur do not matter. For $t\geq 1$, let
$\nu_{t}$ $=$ $\#\left(\{L-1-i_{1},\ldots,L-1-i_{m}\}\cap\{L+t-1-i_{1},\ldots,L+t-1-i_{m}\}\right).$ (22)
@@
where $\wedge$ denotes the bitwise AND operation, and $\gg t$ denotes right shift by $t$ places. Define